Given the following demand equation Qd = 100 - 15P + 20I - 10Pc + 20Ps where I is income and is equal to 200; Pc is the price of a complement and is equal to 30; and Ps is the price of a substitute and is equal to 60. The supply curve is given as Qs = 15P.

Calculate the equilibrium price and quantity, Please explain in detail.
Answer #1

Given Qd=100-15P+20I-10Pc+20Ps

I=200 ; Pc=30;Ps=60

And also given supply curve Qs=15P

Demand curve

Qd=100-15P+20(200)-10(30)+20(60)

=100-15P+4000-300+1200

Qd =5000-15P

At equilibrium

Demand = supply

5000-15P=15P

5000=15P+15P

   30P=5000

P=$166.67

Substitue in Qd=5000-15(196.67)=5000-2500=2500

Therefore equilibrium price =$166.67 and equilibrium quantity=2500